# Mechanical Squeezed-Fock Qubit: Towards Quantum Weak-Force Sensing

### v1

@@ Abstract
We next estimate the fundamental sensitivity through signal-to-noise ratio (SNR) analysis, which is defined as [62]
$\text{SNR}=\frac{\delta P_{|1\rangle_{S}}}{\sigma_{p}}=e^{-\gamma_{0}\cosh(2r)t/2}x_{0}^{2}e^{2r}\delta kC\sqrt{N}t$ (12)
where $\sigma_{p}=1/2C\sqrt{N}$ is the total noise with the number of experiment $N=T/(t+t_{m})$, $T$ is the total available measurement time and $t_{m}$ is the extra time needed to initialize, manipulate, and read out the sensor, $C$ represents the efficiency parameter of readout. We take the value $C=1$ and $t_{m}=0$. The sensitivity is then defined as the minimum detectable signal $\delta k_{\text{min}}$ that yields unit SNR for an integration time of $T=1$[55]. Its minimal over the encoding time is
@@
which is obtained in the condition of $d\delta k_{\text{min}}/dt=0$, i.e., $t=1/[\gamma_{0}\cosh(2r)]$. The sensitivity is exponentially promoted by a factor $\sim e^{r}$ compared to conventional mechanical Fock qubits. We have shown the sensitivity $\delta k_{\text{min}}/k_{0}$ in Fig. 3(c), where $k_{0}=\sqrt{\gamma_{0}e}/x_{0}^{2}$ is the standard sensitivity of $kx^{2}/2$ sensing by the mechanical Fock qubits [76]. Here, The red dashed line indicates the reference $\delta k_{\text{min}}/k_{0}=1$. We can clearly see that $\delta k_{\text{min}}/k_{0}$ decreases as $r$ increases, which indicates that our proposal provides higher precision for quantum sensing.
@@
Experimental feasibility.—To explore the experimental feasibility of our proposal, we here discuss the relevant parameters. We consider a mechanical resonator with frequency $\omega_{a}/2\pi=600$ MHz and decoherence $\gamma_{0}/2\pi=3$ kHz, which results in a quality factor $Q=2\times 10^{5}$. For a finite temperature of 10 mK, the average thermal phonon numbers $\overline{n}\approx 0.00595\ll 1$ and can be ignored. Assuming a weak Kerr nonlinear with strength $K=3$ kHz, we apply a two-phonon drive to the mechanical mode satisfying $\Omega_{p}/\delta_{a}=\tanh(2r)$ with $r=1.5$. In this case, the anharmonicity and decoherence rate are then enhanced to $\sim 1.8$ MHz and $\sim 13.5$ kHz, where we have $\gamma\ll\alpha$, such that the mechanical resonator can be seen as a two-level system with a frequency $\omega_{b}=4.2$ MHz. Mechanical resonators with weak nonlinearity have been reported in nanomechanical systems such as cantilever, membranes or carbon nanotubes [78, 79, 80, 81, 82, 83]. In addition, our proposed scheme is also applicable to the systems whose nonlinearities are introduced by coupling to two energy level systems such as superconducting qubits [84, 85, 40, 52].
@@
We next consider using this mechanical squeezed Fock qubit to sense weak force filed $kx^{2}/2$. For a typical example of carbon nanotube resonators with frequency $\omega_{a}$, the mass is $m=10^{-21}$ kg, we have the zero-point fluctuation $x_{0}=4\times 10^{-12}$ m. Under a decoherence rate of $\gamma_{0}/2\pi=3$ kHz, the calculated sensitivity reaches $\delta k_{\text{min}}=4.71\times 10^{-10}$ $\text{N}\cdot\text{m}^{-1}/\text{Hz}$. This represents an order-of-magnitude improvement over the undriven case with identical parameters ($k_{0}=2.11\times 10^{-9}$ $\text{N}\cdot\text{m}^{-1}/\text{Hz}$), and the enhancement becomes more pronounced at larger values of $r$.
@@
Conclusion.—In conclusion, we have shown that, bt applying a two-phonon drive to mechanical resonators with weak Kerr nonlinearities, one can obtain an exponentially enhanced, tunable anharmonicities. Though the decoherence is also expeonentialy enhanced, the main benefit is that the enhancement of the former is much greater than that of the latter. This scheme enables us to encode a mechanical qubit in the first two squeezed-Fock phonon states. The mechanical qubit can couple to a variety of forces and thus be used to sense weak signals. Compared to the traditional mechanical Fock qubit, the sensitivity using the squeezed-Fock qubit has been exponentially promoted. The proposed squeezed-Fock state encoding is general and can be applied to other bosonic systems like photons and magnons.

### v2

@@ Abstract
We next estimate the fundamental sensitivity through signal-to-noise ratio (SNR) analysis, which is defined as [Degen2017]
$\text{SNR}=\frac{\delta P_{|1\rangle_{S}}}{\sigma_{p}}=e^{-\gamma_{0}\cosh(2r)t/2}x_{0}^{2}e^{2r}\delta kC\sqrt{N}t,$ (12)
where $\sigma_{p}=1/2C\sqrt{N}$ is the total noise with the number of experiment $N=T/(t+t_{m})$, $T$ is the total available measurement time and $t_{m}$ is the extra time needed to initialize, manipulate, and read out the sensor, $C$ represents the efficiency parameter of readout. We take the value $C=1$ and $t_{m}=0$. The sensitivity is then defined as the minimum detectable signal $\delta k_{\text{min}}$ that yields unit SNR for an integration time of T=1T=1[Pistolesi2021]. Its minimal over the encoding time is
@@
which is obtained in the condition of $d\delta k_{\text{min}}/dt=0$, i.e., $t=1/[\gamma_{0}\cosh(2r)]$. We have shown the relative sensitivity $\delta k_{\text{min}}/\delta k_{0}$ in Fig. 3(c), where $\delta k_{0}=\sqrt{\gamma_{0}e}/x_{0}^{2}$ is the standard sensitivity of $kx^{2}/2$ sensing by the mechanical Fock qubit [SM]. Here, the red dashed line indicates the reference $\delta k_{\text{min}}/\delta k_{0}=1$. We can clearly see that $\delta k_{\text{min}}/\delta k_{0}$ decreases as $r$ increases. The sensitivity is exponentially promoted by a factor $\sim e^{r}$ compared to conventional mechanical Fock qubits, which indicates that our proposal provides higher precision for quantum sensing.
@@
Experimental feasibility.—To explore the experimental feasibility of our proposal, we here discuss the relevant parameters. We consider a mechanical resonator with frequency $\omega_{a}/2\pi=600$ MHz and decoherence $\gamma_{0}/2\pi=3$ kHz. For a finite temperature of 10 mK, the average thermal phonon numbers $\overline{n}\approx 0.00595\ll 1$ and can be ignored. Assuming a weak Kerr nonlinear strength $K/2\pi=3$ kHz, we apply a two-phonon drive to the mechanical mode satisfying $\Omega_{p}/\delta_{a}=\tanh(2r)$ with $r=1.5$. In this case, the anharmonicity and decoherence rate are then enhanced to $\sim 1.8$ MHz and $\sim 13.5$ kHz, where we have $\Gamma\ll\alpha$, such that the mechanical resonator can be seen as a two-level system with a frequency $\omega_{b}/2\pi=4.2$ MHz. Mechanical resonators with weak nonlinearity have been reported in nanomechanical systems such as cantilevers, membranes or carbon nanotubes [Kim2013, Djorwe2014, ArellanoCastro2014, Ella2015, Lai2022, Venkatachalam2023]. In addition, our proposed scheme is also applicable to the systems whose nonlinearities are introduced by coupling to two energy level systems such as superconducting qubits [Liu2010, Ramos2013, Wollman2015, Chu2017].
@@
We next consider using this mechanical squeezed Fock qubit to sense weak force field $kx^{2}/2$. For a typical example of carbon nanotube resonators with frequency $\omega_{a}$, the mass is $m=10^{-21}$ kg, we have the zero-point fluctuation $x_{0}=4\times 10^{-12}$ m. Under a decoherence rate of $\gamma_{0}/2\pi=3$ kHz, the calculated sensitivity reaches $\delta k_{\text{min}}=4.71\times 10^{-10}$ $\text{N}\cdot\text{m}^{-1}/\text{Hz}^{1/2}$. This represents an order-of-magnitude improvement over the undriven case with identical parameters ($\delta k_{0}=2.11\times 10^{-9}$ $\text{N}\cdot\text{m}^{-1}/\text{Hz}^{1/2}$), and the enhancement becomes more pronounced at larger values of $r$.
@@
Conclusion.—In conclusion, we have shown that, by applying a two-phonon drive to mechanical resonators with weak Kerr nonlinearities, one can obtain an exponentially enhanced and tunable anharmonicity. Though the decoherence is also expeonentialy enhanced, the main benefit is that the enhancement of the former is much greater than that of the latter. This scheme enables us to encode a mechanical qubit in the first two squeezed-Fock phonon states. The mechanical qubit can couple to a variety of forces and thus be used to sense weak signals. Compared to the traditional mechanical Fock qubit, the sensitivity using the squeezed-Fock qubit has been exponentially promoted. The proposed squeezed-Fock state encoding is general and can be applied to other bosonic systems like photons and magnons.